thatlightskinting:

One of the games in a fundraiser involves rolling a six-sided die. The rule is that a player has to donate six times the value shown on the die. What is the expected value of a donation? A. $21 B. $27 C. $30 D. $36

Vocaloid:

 expected value (EV) formula above. It's a lot simpler than it looks. for every outcome, multiply: (the value of that outcome)(the probability of getting that outcome) then, once you've calculated this for all outcomes, add the products together.

Vocaloid:

I'll walk you through part of the problem. if you roll a 1 on the die, that means you have to donate $6 (since 1 * 6 = 6). the probability of getting 1 on a die is 1/6. so the expected value so far is 6 * (1/6) = 1 if you roll a 2 on the die, that means you have to donate $12 (since 2 * 6 = 12). the probability of getting 2 on a die is 1/6. so the expected value for this outcome is 12*(1/6) = 2 repeat for all 6 outcomes, then add up the results